# The Z-Gromov-Wasserstein Distance

### v1

@@ 5.2 Approximation of $Z$-network GW distances via $\mathbb{R}^{n}$-networks
This can be understood via the Frechet-Kuratowski theorem [55, Proposition 1.17], which isometrically embeds any separable metric spaces to the $\ell^{\infty}$ space. The embedding $Z\hookrightarrow\ell^{\infty}$ for bounded $Z$ is provided by $z\mapsto\{d_{Z}(z,z_{i})\}_{i=1}^{\infty}$ where $\{z_{i}\}_{i=1}^{\infty}$ is the dense subset of $Z$, so the mapping $\omega_{X}\mapsto\omega_{X_{Q}}=\{d_{Z}(\omega_{X},q_{i})\}_{i=1}^{n}$ works as an approximation of the embedding, and this is why the GW distances are equal in the limit of $Q\to Z$. We note that although the inequality is true with just the boundedness assumption, it is necessary to impose a stronger condition, such as compactness, to make the Hausdorff distance arbitrarily small. In the compact case, previous works such as [12, Theorem 34], [16, Theorem 2] show bounds on the convergence rate to 0 of the Hausdorff distance $\textnormal{H}(Z,Q)$ as $n\to\infty$ when $Q=\{q_{1},\cdots q_{n}\}$ is a random sample from $Z$.

### v2

@@ 5.2 Approximation of $Z$-network GW distances via $\mathbb{R}^{n}$-networks
This can be understood via the Frechet-Kuratowski theorem [64, Proposition 1.17], which isometrically embeds any separable metric spaces to the $\ell^{\infty}$ space. The embedding $Z\hookrightarrow\ell^{\infty}$ for bounded $Z$ is provided by $z\mapsto\{d_{Z}(z,z_{i})\}_{i=1}^{\infty}$ where $\{z_{i}\}_{i=1}^{\infty}$ is the dense subset of $Z$, so the mapping $\omega_{X}\mapsto\omega_{X_{Q}}=\{d_{Z}(\omega_{X},q_{i})\}_{i=1}^{n}$ works as an approximation of the embedding, and this is why the GW distances are equal in the limit of $Q\to Z$. We note that although the inequality is true with just the boundedness assumption, it is necessary to impose a stronger condition, such as compactness, to make the Hausdorff distance arbitrarily small. In the compact case, previous works such as [14, Theorem 34], [18, Theorem 2] show bounds on the convergence rate to 0 of the Hausdorff distance $\textnormal{H}(Z,Q)$ as $n\to\infty$ when $Q=\{q_{1},\cdots q_{n}\}$ is a random sample from $Z$.
@@
6 Discussion
The $Z$-Gromov-Wasserstein framework defines a very general setting for reasoning about mathematical properties of GW-like distances which have appeared in the recent literature. These properties have frequently been derived independently in each instance; in contrast, the work here provides a high-level perspective, showing that the metric properties of a $Z$-GW distance are closely related to those of the space $Z$. This initial paper on the topic is meant to lay out its foundations, and we envision several directions for future research, both theoretical and applications-oriented in nature. We conclude the paper by outlining some important future goals:
Geodesic structure and curvature bounds. The geodesic structure of the $Z$-GW space $\mathcal{M}_{\sim}^{Z,p}$, and how it relates to that of $Z$, has so far only been partially characterized in Theorem 7. Answers to the related open questions, 3 and 4, would provide a more complete characterization. Moreover, it would be interesting to give conditions under which the geodesics of $Z$-GW distance are always of the form described explicitly in the proof of Theorem 7. For example, a result of this form is demonstrated by Sturm for standard GW distance in [77], and this is generalized to certain GW variants in [84, 26, 91]. Such a characterization would be a first step in establishing bounds on the Alexandrov curvature of the $Z$-GW space. Understanding the relationship between the curvature of the space $Z$ and that of the $Z$-GW space would be very interesting from a theoretical perspective, and could have practical implications to algorithm design. For example, curvature estimates can be used to inform algorithms for computing Fréchet means of point clouds in a metric space—see, e.g., [82, 24, 91] for the case of lower bounds, or [31, 3] for the case of upper bounds.
Topological Questions. We showed in Theorem 6 that, when $p<\infty$, the $Z$-GW space is contractible, perhaps surprisingly, independently of the topology of $Z$. However, the dependence of the topology of $\mathcal{M}_{\sim}^{Z,\infty}$ on that of $Z$ has not been resolved—see 1 and 2. We show in Example 4.12 and the ensuing discussion that if $Z$ is not path-connected, then neither is $\mathcal{M}_{\sim}^{Z,\infty}$, so the answer to these questions is potentially subtle. We remark here that the special case of $p=\infty$ is conceptually important—the distance $\mathrm{GW}_{\infty}^{Z}$ is essentially a notion of Gromov-Hausdorff distance for $Z$-valued networks, and Gromov-Hausdorff distance for $\mathbb{R}$-valued networks is already a rich area of study, with ties to areas such as topological data analysis [20, 21]. Other interesting topological questions remain open. For example, Gromov’s celebrated precompactness theorem (see [13, Chapter 7]) says that a family of metric spaces with certain uniform bounds on geometrical properties of its elements has compact closure in the Gromov-Hausdorff topology, and this is generalized to the Gromov-Wasserstein setting in [62, Theorem 5.3]. A natural question is whether there are interesting precompact families for $Z$-networks in the $Z$-GW topology.
Structure of Optimal Couplings. There has recently been significant interest in characterizing the structure of optimal couplings in the GW framework; in particular, several recent articles address the problem of finding general classes of measure networks under which the GW distance is realized by a measure-preserving map (e.g. [57, 28, 77, 30, 27]). The $Z$-GW framework gives a more general setting for addressing these questions. The added flexibility in the model might allow for more tractable versions of the problem to be attacked.
Computational Pipeline. The focus of this paper is on the basic theory of $Z$-GW distances, but the results in Section 5 point toward more practical considerations. The lower bounds in Theorem 8 are polynomial-time computable in the standard GW setting; the computational efficiency for some of these lower bounds relies on the fact that Wasserstein distances between distributions on the real line can be computed via an explicit formula. From an applications perspective, it would be useful to characterize other spaces $Z$ where the Wasserstein distance is efficiently and explicitly computable (e.g., the circle [66]), and to utilize this for efficient $Z$-GW distance lower bound computation. Another interesting line of research in this direction is to determine classes of $Z$-networks for which the lower bounds are injective in the sense that vanishing of the lower bound implies that the input $Z$-networks are weakly isomorphic—this question is quite subtle in the standard GW setting [58]. Finally, an important consquequence of Theorem 9 is that, for arbitrary $Z$, $Z$-GW distances can be approximated via solvers for the $\mathbb{R}^{n}$-GW distance. Algorithms for approximating the latter distance should follow by straightforwardly adapting solvers for the standard GW distance. A serious implementation of this approximation scheme is an important next step.
Applications. The surplus of examples of $Z$-GW distances provided in Section 3 suggests the wide applicability of this framework. Once the computational pipeline described in the previous paragraph is in place, we plan to apply it to various problems involving analysis of complex and non-standard data. We are particularly interested in exploring the more novel settings described in Section 3.2, such as probabilistic metric spaces, shape graphs and connection graphs.